OpenStudy (anonymous):

A rectangular garden 200 square feet in area is to be fenced off against rabbits. Find the dimensions that will require the least amount of fencing if one side of the garden is already protected by a barn.

OpenStudy (anonymous):

Your question is unclear. Is soemthing missing?

OpenStudy (anonymous):

no that's the entire question in the book :/

OpenStudy (mathstudent55):

 A = xy Do you have any information about how much fencing material there is?

OpenStudy (anonymous):

nope :/

OpenStudy (mathstudent55):

Length of fence = 2x + y Area = xy = 200 ----> y = 200/x Fence: 2x + y = 2x + 200/x F = 2x + 200/x F' = 2 - 200/x^2 2 - 200/x^2 = 0 -200/x^2 = -2 x^2 = 400 \(x = \sqrt{400} \) x = 20 A = xy = 200 20y = 200 y = 10

OpenStudy (anonymous):

thank you very much :)

OpenStudy (mathstudent55):

wlcm